a) Mr. David borrowed RM125,000 to buy a car. His loan cost was 12% and he promised to repay the loan in 15 equal annual payments. How much is the principal outstanding after the first loan payment?



b) Puan Elly is planning to retire in 20 years, and she plans to begin withdrawing RM50,000 annually at that point of time. She is expected to live for 20 years following her retirement. If she can earn 9% compounded annually, how much does she need to save each year to prepare for her financial needs after her retirement? (Total 10 marks)



c) Rafael Corporation expects to receive RM20,000 per year for 10 years and RM35,000 per year for the next 10 years. What is the present value of this cash flow given 12% discount rate? (Total 5 marks)

 





1
Expert's answer
2022-03-09T03:02:03-0500

a. 140,000 - 9333.33 = 130666.66

b. (91*50000)/100 = RM45500

c. (88*35000)/100 = RM30800